OpenStudy (anonymous):

Solve the equation by factoring: 4x^2=25

OpenStudy (campbell_st):

well rewrite the equation and you get \[4x^2 - 25 = 0\] this is the difference of 2 squares.... which can be factored... hope it helps

OpenStudy (campbell_st):

then when you have to factors... let each factor equal zero... and then solve for x

OpenStudy (anonymous):

Ok i get how that one works but what about: x^2+2x=y

OpenStudy (campbell_st):

well its a parabola... which can be factored as y = x(x + 2) so it cuts the x - axis at x = 0 and x = - 2 the parabola is concave up and has a minumum value when x = -1 so the minimum value is y = -1

OpenStudy (anonymous):

Thanks for all the help!

OpenStudy (campbell_st):

glad to help good luck